. Sheryl left by plane to visit her mother in Hartford, 42 km away. Fifteen minutes later, her mother left to meet her at the airport. She drove the 20 km to the airport at 40 km per hr, arriving just as the plane taxied in. What was the speed of the plane?

Answer by stanbon(75887)

Sheryl left by plane to visit her mother in Hartford, 42 km away.
Fifteen minutes later, her mother left to meet her at the airport.
She drove the 20 km to the airport at 40 km per hr, arriving just as the plane taxied in.
What was the speed of the plane?
---------------------
Mom's DATA:
Distance = 20 km; rate = 40 kmh ; time = d/r = 1/2 hr
---------------------
Plane DATA;
Distance = 42 km ; time = (1/4 hr)+(1/2 hr) = (3/4 hr)
---------------
EQUATION:
Plane Rate = 42km/(3/4)hr = 56 kph
====================
Cheers,
Stan H.